Prepare the cash flow statement from investing activities of Alpha Creative Ltd for the year ended March31, 2019 -

Particulars Amount

Plant acquired 160000

Claim received for loss of plant in fire 45500

Unsecured loans given to subsidiaries 595000

Interest on loan received from subsidiary companies 72500



• Also give reasons for the classification of above activities as inflow /outflow
1
Expert's answer
2019-11-25T09:40:28-0500

Cash Flow from Investing Activities is the section of a company’s cash flow statement that displays how much money has been used in (or generated from) making investments during a specific time period. Investing activities include purchases of long-term assets (such as property, plant, and equipment), acquisitions of other businesses, and investments in marketable securities (stocks and bonds).

Cash receipts: 45500 + 72500 = 118000.

Cash paid: 160000 + 595000 = (755000).

Net cash flow: 118000 - 755000 = (637000).